# Funding AI for Good: A Call for Meaningful Engagement

### v2

@@ Appendix A Funding documents
See Table LABEL:tab:funding_doc_mr next page.
@@
C12 Innovate Africa Challenge Artificial Intelligence (AI) for Climate Action 2024 \$50,000 multi-phased (bootcamp; incubation; implementation) BMZ Government Agency Germany Africa climate change “company from any of the 39 Smart Africa Alliance countries” (Smart Africa, 2024)
C13 Call for Proposals: Climate Change AI Innovation Grants 2024 up to \$150,000 per proposal (up to \$1.4M total) 1 year Climate Change AI (QCF, Google DeepMind, Global Methane Hub) Non-profit United States Global climate change “PI affiliated with accredited university in one of 38 OECD member countries” (Climate Change AI, 2024)
C14 Request for Proposals: Natural Language Processing (NLP) 2024 \$1,000,000 1.5 years Lacuna Fund (supported by Google.org) Non-profit United States Africa and Latin America NLP data “non-profit entity, research institution, for-profit social enterprise, or team of such orgs” (Lacuna Fund, 2024a)
C15 Request for Proposals: Antimicrobial Resistance (AMR) 2024 \$1,000,000 1.5 years Lacuna Fund (supported by Wellcome Trust) Non-profit United Kingdom LMICs in Africa, Latin America, South & Southeast Asia public health “non-profit entity, research institution, for-profit social enterprise, or team of such orgs” (Lacuna Fund, 2024b)
@@
C17 Big Data Science: Mechanisms of HIV Activity & Interaction with Substance Use Disorder 2023 \$2,000,000 Up to 5 years NIH Government Agency United States Global public health Organizations (National Institutes of Health (2025), NIH)
@@
C19 Smart Health & Biomedical Research in the Era of AI & Advanced Data Science (SCH) 2025 Up to \$20,000,000 per year Up to 4 years NSF Government Agency United States United States public health “Institutions of Higher Education; Non-profit, non-academic organizations” (U.S. National Science Foundation, 2025)
A1 “Patrick J. McGovern Foundation commits \$40M to advancing AI and data for good” 2021 \$40,000,000 not specified Patrick J. McGovern Foundation Philanthropy Foundation United States unspecified broadly nonprofits “Nonprofits and NGOs; orgs on data dignity, data stewardship, AI ethics” (Foundation, 2021)
A2 “Artificial Intelligence to Accelerate Inclusion Challenge” 2024 not specified not specified Mastercard Center for Inclusive Growth; data.org Non-profit United States Global economic empowerment “the diverse range of communities often left out of AI for social impact” (Mastercard, 2024)

### v3

@@ Appendix A Funding documents
See Table LABEL:tab:funding_doc_mr next page. The full dataset can be accessed through this link:
https://github.com/nicole-hjlin/chi2026-ai-for-good-funding-dataset.
@@
C12 Innovate Africa Challenge Artificial Intelligence (AI) for Climate Action 2024 \$50,000 multi-phased (bootcamp; incubation; implementation) BMZ Government Agency Germany Africa climate change company from any of the 39 Smart Africa Alliance countries (Smart Africa, 2024)
C13 Call for Proposals: Climate Change AI Innovation Grants 2024 up to \$150,000 per proposal (up to \$1.4M total) 1 year Climate Change AI (QCF, Google DeepMind, Global Methane Hub) Non-profit United States Global climate change PI affiliated with accredited university in one of 38 OECD member countries (Climate Change AI, 2024)
C14 Request for Proposals: Natural Language Processing (NLP) 2024 \$1,000,000 1.5 years Lacuna Fund (supported by Google.org) Non-profit United States Africa and Latin America NLP data “non-profit entity, research institution, for-profit social enterprise, or team of such orgs” (Lacuna Fund, 2024b)
C15 Request for Proposals: Antimicrobial Resistance (AMR) 2024 \$1,000,000 1.5 years Lacuna Fund (supported by Wellcome Trust) Non-profit United Kingdom LMICs in Africa, Latin America, South & Southeast Asia public health “non-profit entity, research institution, for-profit social enterprise, or team of such orgs” (Lacuna Fund, 2024a)
@@
C17 Leveraging Big Data Science to Elucidate the Mechanisms of HIV Activity and Interaction with Substance Use Disorder 2023 \$2,000,000 Up to 5 years NIH Government Agency United States Global public health Organizations (National Institutes of Health (2025), NIH)
@@
C19 Smart Health & Biomedical Research in the Era of AI & Advanced Data Science (SCH) 2025 Up to \$20,000,000 per year Up to 4 years NSF Government Agency United States United States public health Institutions of Higher Education; Non-profit, non-academic organizations (U.S. National Science Foundation, 2025)
A1 “[Patrick J. McGovern] Foundation commits \$40M to advancing AI and data for good” 2021 \$40,000,000 not specified Patrick J. McGovern Foundation Philanthropy Foundation United States unspecified broadly nonprofits Nonprofits and NGOs; orgs on data dignity, data stewardship, AI ethics (Foundation, 2021)
A2 “Mastercard Center for Inclusive Growth and data.org launch Artificial Intelligence to Accelerate Inclusion challenge” 2024 not specified not specified Mastercard Center for Inclusive Growth; data.org Non-profit United States Global economic empowerment “the diverse range of communities often left out of AI for social impact” (Mastercard, 2024)
@@ Appendix B Coding Process and Affinity Diagram
This figure shows a screenshot of an online whiteboard where sticky notes of emerging themes are grouped into clusters. Arrows are used to link clusters and indicate relationships between them.
@@
This figure shows a screenshot of an online whiteboard where sticky notes of codes are grouped into clusters under different themes. These themes are placed along a spectrum of more balanced and more techno-centric under each funding document component: background (on the top left corner), expected outcomes (on the bottom left corner), eligibility and team composition (on the top right corner), and funders’ support (on the bottom right corner).